OpenStudy (anonymous):

Okay! Solve each trig equation below, make sure to give all possible solutions. (all angles are in degrees.) NO CALCULATOR.

OpenStudy (anonymous):

\[1.) \sin \theta=-1\]

OpenStudy (anonymous):

\[2.) 2\cos \alpha=\sqrt{3}\]

OpenStudy (anonymous):

1.) is -90+n360